Suppose that P(A/B)=0.2, P(A/B^c)= 0.3 and P(B)=0.8 what is P(A) ?
1
Expert's answer
2021-09-21T11:45:36-0400

"P(A)=P(A|B)+P(A|B^c)=0.2+0.3=0.5"